OpenStudy (anonymous):

Set up the following application problem and don't forget to declare your variable. State your solution and round to 2 decimal places. Here's the question: Danielle like rowing her canoe up and down the Chicopee River. She can row her canoe 5 miles downstream in the same time that it takes to row 3 miles upstream. If the current of the river is approximately 2mph, find Danielle's rowing speed in still water

OpenStudy (matt101):

Let her rowing speed be x (speed of water current would be added or subtracted from it) First, remember the equation v = d/t --> t = d/v Since the time it takes to go 5 miles downstream is the same as the time it takes to go 3 miles upstream: 5/(x+2) = 3/(x-2) Solve for x :)